OpenStudy (anonymous):

Can someone tell me how to solve this rational expression please x² - 9 x + 3

OpenStudy (kamille):

x² - 9 x + 3= x² - 9 x + 3=0 D=(-9)^2-4*3=81-12=69 x1=(9+/sqrt69)/2 x2=(9-/sqrt69)/2 the answer x² - 9 x + 3=(x-(9+/sqrt69)/2)((9+/sqrt69)/2)

OpenStudy (anonymous):

Or use the formula which I forgot it's name... I think it's called the quadratic formula. a := 1 b := -9 c := 3 D = Discriminant := b^2 - 4ac you will get a positive D